OpenStudy (anonymous):

find the sum of the first 30 terms of the sequence below plz and explain. A[n]=4[n]+1

jimthompson5910 (jim_thompson5910):

Similar Example: find the sum of the first 18 terms of the sequence A[n] = 7n + 3 Step 1) Find the first term. To do that, we plug in n = 1 A[n] = 7n + 3 A[1] = 7(1) + 3 A[1] = 10 The first term is a1 = 10 ---------------------------- Step 2) Find the last term being added. In this example, the last term is the 18th term and that happens when n = 18 A[n] = 7n + 3 A[18] = 7(18) + 3 A[18] = 129 The last term being added up is a18 = 129 ---------------------------- Step 3) Now use this formula \[\Large S_n = \frac{n(a_1 + a_n)}{2}\] \[\Large S_{18} = \frac{18(a_1 + a_{18})}{2}\] \[\Large S_{18} = \frac{18(10+129)}{2}\] \[\Large S_{18} = \frac{18(139)}{2}\] \[\Large S_{18} = 9(139)\] \[\Large S_{18} = 1,251\] So in my example, if you add up the first 18 terms (a1 through a18) for the sequence A[n] = 7n+3, you'll get the result 1251 Keep in mind that this is just an example and not the solution to your specific problem; however, it's very similar to what you have. So hopefully this helps you out. If not, then tell me where you are getting stuck.

OpenStudy (anonymous):

ty